# A subset $S$ of vertices in a planar graph $G$ is a free set if, for every set $P$ of $|S|$ points in the plane, there exists a straight-line crossing-free drawing of $G$ in which vertices of $S$ are mapped to distinct points in $P$. In this survey, we review

### v1

@@ 4 Applications
In this section we discuss applications of free sets. For each of these ($s$) applications, and each of the ($t$) classes of planar graphs known to have large free sets, the use of free sets gives a result. Rather than provide an exhaustive list of $s\times t$ corollaries, we state the relationship between free sets and the application and give the most general corollary; the one about planar graphs obtained from 5.
@@ 4.1 Untangling
Let $\mathcal{G}$ be the class of all planar graphs. For every $n$-vertex graph $G\in{\mathcal{G}}$, $\operatorname{fix}(G)\in\Omega(n^{1/4})$. In other words, $\operatorname{fix}_{\mathcal{G}}(n)\geq(n/2)^{1/4}$.
@@
On the contrary, applying the known bounds on the largest free sets coupled with Lemma 1 gives tight bounds for some graph families. Specifically, the linear bounds on free sets in Section 3.1 and Lemma 1 imply $\Omega(n^{1/2})$ untangling bound for these families. Bose et al. [17] showed that there is a class of trees, in fact forest of stars, ${\mathcal{T}}$, such that $\operatorname{fix}_{\mathcal{T}}\in O(n^{1/2})$. Thus the $\Omega(n^{1/2})$ untangling bound is tight for all families in Section 3.1, except for the class of cubic triconnected planar graphs. The stars in ${\mathcal{T}}$ have unbounded degree. Thus a tight bound on untangling cubic triconnected planar graph is unknown.
@@ 4.2 Universal point subsets
A set of points $P$ is universal for a set of planar graphs if every graph from the set has a straight-line crossing-free drawing where all of its vertices map to distinct points in $P$. Arguably, the most famous open problem in graph drawing, attributed to Bojan Mohar (1988), asks if there exist a constant $c$ such that, for every $n$ there exists a pointset of size $c\cdot n$ that is universal for the class of all $n$-vertex planar graphs. It is known that for all, large enough $n$, no universal pointset of size $n$ exists for the class of $n$-vertex planar graphs – as first proved by de Fraysseix et al. [34]. The same authors also proved that the $O(n)\times O(n)$ integer grid is universal for all $n$-vertex planar graphs and thus universal pointsets of size $O(n^{2})$ exists. Currently the best known lower bound on the size of a smallest universal pointset for $n$-vertex planar graphs is $1.239n-o(n)$ [50] and the best known upper bound is $n^{2}/4-O(n)$ [6]. Closing the gap between $\Omega(n)$ and $O(n^{2})$ is a major, and likely very difficult, graph drawing problem, open since 1988 [33, 34].
Various related notions have been introduced and studied in the literature with the aim of better understanding the universal pointset problem. A set $P$ of $k$ points in the plane is a universal point subset for a family of planar graphs, if for every graph $G$ in the family, there exists a straight-line crossing-free drawing in which $k$ vertices of $G$ are placed at the $k$ points in $P$. Universal point subsets were introduced by Angelini et al. [2]. They proved that (a very flat convex chain of) $\lceil\sqrt{n}\rceil$ points is a universal point subset for the class of all $n$-vertex planar graphs. Giacomo et al. [37] then proved that every set of $\Omega(\log n)$ points, as well as, every set of $\lceil n^{1/3}\rceil$ points in convex position is a universal point subset for the class of $n$-vertex planar graphs.
@@ 4.3 Simultaneous Geometric Embeddings
Simultaneous geometric embeddings were introduced by Braß et al. [19]. Since then there has been a plethora of work on the subject on many variants of the problem– see, for example a survey by Bläsius et al. [13]. Common variants of the problem include those in which the mapping between the vertices of the two graphs is given and those in which the mapping is not given.
@@ 4.3.1 Without Mapping
A sequence of graphs $G_{1},G_{2},\dots,G_{r}$ where $|V(G_{1})|\geq|V(G_{2})|\geq\dots|V(G_{r})|$ are said to have a simultaneous geometric embedding without mapping (SGE-nomap) if there exists a pointset $P$ of size $|V(G_{1})|$ such that each of $G_{1},\ldots,G_{r}$ has a straight-line crossing-free drawing where all of its vertices are mapped to distinct points in $P$. The following well known and still wildly open problem was asked by Braß et al. [19] in 2003, and is also listed among selected list of graph drawing problems in [18] (see also Problem 12): Does every pair of $n$-vertex planar graphs has SGE-nomap (for every positive integer $n$)? The statement is known not to be true when “pair” (that is, $r=2$) is replaced by a bigger constant [21, 51], currently being $r=30$ [54]. More generally, Steiner [54] showed that for every large enough $n$ there exists a sequence of at most $(3+o(1))\log_{2}(n)$ $n$-vertex planar graphs that do not have a SGE-nomap.
While the problem of Braß et al. [19] still seems to be out of reach, in a different direction Angelini et al. [3] write: “What is the largest $k\leq n$ such that every $n$-vertex planar graph and every $k$-vertex planar graph admit a geometric simultaneous embedding with no mapping? Surprisingly, we are not aware of any super-constant lower bound for the value of $k$?” The corollary below, noticed first by Dujmović [27], answered their question with the help of free sets, as detailed in the next lemma.
@@
Let $G_{1}$ be a planar graph with free set of size at least $k$. Let $G_{2}$ be a planar graph on at most $k$ vertices. Then $G_{1}$ and $G_{2}$ admit SGE-nomap.
@@
One can combine the many lower bounds (on the size of the largest free set in a graph family) obtained in Sections 3.1, 3.2 and 3.3 to obtain results of similar flavour to Corollary 7. One example is the following, for every $n$ and every $k\leq n/4$, every $n$-vertex planar graph and every $k$-vertex triconnected cubic planar graph admit a SGE-nomap.
Note that universal point subsets of size $k$ do not imply results on SGE-nomap, akin to 7. In particular, imagine that, as in the proof of Lemma 10, one starts with a straight-line crossing-free drawing of the smaller graph $G_{2}$. Even if the larger graph $G_{1}$ is in a class of graphs that have universal point subsets of size $k$, there is no guarantee that $G_{1}$ can be drawn on the specific pointset of size $k$ used by the drawing of $G_{2}$. To be able to use universal point subsets for SGE-nomap problems, one needs the stronger variant of “all pointsets of size $k$” to be subset universal—something that is guaranteed by free sets of size $k$.
@@ 4.3.2 With Mapping
Two $n$-vertex planar graphs $G_{1}$ and $G_{2}$ on the same vertex set, $V:=V(G_{1})=V(G_{2})$, are said to have a $k$-partial simultaneous geometric embedding with mapping ($k$-PSGE-withmap) if there exists a set $V^{\prime}:=\{v_{1},\ldots,v_{k}\}\subseteq V$, and a set $P:=\{p_{1},\ldots,p_{k}\}$ of points such that each of $G_{1}$ and $G_{2}$ has a straight-line crossing-free drawing in which $v_{i}$ is mapped to $p_{i}$, for each $i\in\{1,\ldots,k\}$. PSGE-withmap on the whole vertex set (i.e. $n$-PSGE-withmap) has been widely studied leading to mostly negative results (thus giving another motivation to introduce this partial version). For example, it is known that, for every large enough $n$, there are pairs of $n$-vertex planar graphs that do not have $n$-PSGE-withmap [19]. In fact the same is true for graphs from very simple families of planar graphs, for example: for an $n$-vertex tree and an $n$-vertex path [5], for an $n$-vertex planar graph and an $n$-vertex matching [5] and for three $n$-vertex paths [19].
The $k$-PSGE-withmap problem was introduced [^8] In [31], they use the abbreviation $k$-PSGE for what we call $k$-PSGE-withmap. Also, the definition of $k$-PSGE-withmap, as introduced in Evans et al. [31], has one additional requirement, which states that if $v,w\in V$ are mapped to the same point in $D_{i}$ and $D_{j}$, then $v=w$. However this additional requirement can always be met by the fact that it is possible to perturb any subset of vertices in a straight-line crossing-free drawing without introducing crossings. (This fact is used for example in the proof that every free-collinear set is free.) by Evans et al. [31] who proved that any two $n$-vertex trees have an $1711n/$ 17-PSGE-withmap. Their proof uses their column planarity, which is the topic of the next section. Barba et al. [9] proved that any two $n$-vertex outerplanar graphs have an $n/4$-PSGE-withmap. Evans et al. [31] also observed that 5, the untangling result, implies that every pair of $n$-vertex planar graphs has an $\Omega(n^{1/4})$-PSGE-withmap. Namely, start with a straight-line crossing-free drawing of $G_{1}$. Since $G_{1}$ and $G_{2}$ have the same vertex set, the drawing of $G_{1}$ (or rather the positions of its vertices in the plane) defines a straight-line drawing of $G_{2}$ (that almost certainly has crossings). Untangling $G_{2}$ while keeping $\Omega(n^{1/4})$ of its vertices fixed (which is possible by 5) gives the result.
@@
However, the above untangling argument fails if we try to apply it one more time. The following generalization of the $k$-PSGE problem (to more than two graphs) illustrates this. Given any set $Q:=\{G_{1},\ldots,G_{r}\}$ of planar graphs on the same vertex set, $V$, we say that the graphs in $Q$ have a $k$-partial simultaneous geometric embedding with mapping ($k$-PSGE-withmap) if there exists a set $V^{\prime}:=\{v_{1},\ldots,v_{k}\}\subseteq V$, and a set $P:=\{p_{1},\ldots,p_{k}\}$ of points such that each graph in $Q$ has a straight-line crossing-free drawing in which $v_{i}$ is mapped to $p_{i}$, for each $i\in\{1,\ldots,k\}$.
@@
We start by presenting a lemma about PSGE-withmap for two graphs via free-sets. Its proof uses the common trick of taking advantage of the fact that points in the plane have two degrees of freedom, and thus one ordering can be imposed on the ${x}$-coordinates and a different ordering on ${y}$-coordinates (see for example [8]).
@@
6 and 11 gives another proof of 7. By 6 (applied to $G_{1}$ and $V$), there exists $S_{1}\subseteq V$ of size at least $\sqrt{n/2}$ that is a free set in $G_{1}$. By 6 (applied to $G_{2}$ and $S_{1}$), there exists $S\subseteq S_{1}$ of size at least $\sqrt{|S_{1}|/2}\geq n^{1/4}/2^{3/4}$. Thus $S$ is a free set in $G_{1}$ and in $G_{2}$. By Lemma 11, $G_{1}$ and $G_{2}$ have a $\lceil n^{1/4}/2^{3/4}\rceil$-PSGE-withmap. In exactly the same way, Lemmata 3 and 11 give a proof that any two $n$-vertex graphs each of which is one of the following: a tree, an outerplanar graph, Halin graph, or a square graph, have a $(n/4)$-PSGE-withmap.
@@
Now apply Lemma 11 to the graphs $G_{1}$ and $G_{2}$ with the set $S$ to obtain straight-line crossing-free drawing of both graphs in which each vertex in $S$ (and therefore, the vertices in $S^{\prime}$) appear at the same location in both drawings. In these drawings, the ${y}$-coordinates of the vertices in $S^{\prime}$ are increasing. By the definition of free set (exchanging the roles of ${x}$- and ${y}$-coordinates), each of $G_{3},\ldots,G_{k}$ has a straight-line crossing-free drawing  in which each vertex in $S^{\prime}$ has the same location in each drawing. ∎
@@
If the graphs in $Q$ each come from one of the classes covered by 1, then repeated applications of 1 gives a set $S$ of size at least $n/2^{r}$ that is free in each of $G_{1},\ldots,G_{r}$, in which case the bound in 8 becomes $n^{1/2^{r-2}}/4$.
@@ 4.4 Column planarity
Given a planar graph $G$, a set $R\subseteq V(G)$ is column planar in $G$ if the vertices of $R$ can be assigned distinct ${x}$-coordinates such that for any arbitrary assignment of ${y}$-coordinates to the vertices in $R$ (under the condition that the resulting $|R|$ points have no three on a line), there exists a straight-line crossing-free drawing of $G$ in which each vertex $v\in R$ is placed to its assigned coordinates. $R$ is strongly column planar if the “no three on a line condition” is removed from the definition. The following lemma is immediate from the definition of free sets and strongly column planar sets.
@@
A (sub)set of vertices of a planar graph is strongly column planar set if and only if it is an unordered free set.
@@
Column planar sets were introduced by Evans et al. [31] motivated by applications to partial simultaneous geometric embeddings. Notions similar to column planarity were studied by Estrella-Balderrama et al. [30] and Di Giacomo et al. [25]. Barba et al. [8] proved that $n$-vertex trees have column planar sets of size $1714n/$ 17. Note that that does not imply that trees have free sets of that size, since being column planar does not imply being strongly column planar. In particular, if $G$ is a 3-cycles, then $V(G)$ is column planar set for any assignment of ${x}$-coordinates to $V(G)$. However, no ${x}$-coordinate assignment can be turned into a straight-line crossing-free drawing with all the vertices having the same ${y}$-coordinate. Thus $V(G)$ is not strongly column planar.

### v2

@@ 4 Applications
In this section, we discuss applications of free sets. For each of these ($s$) applications, and each of the ($t$) classes of planar graphs known to have large free sets, the use of free sets gives a result. Rather than providing an exhaustive list of $s\times t$ corollaries, we state the relationship between free sets and the application and give the most general corollary, the one about planar graphs obtained from 5.
@@ 4.1 Untangling
Let $\mathcal{G}$ be the class of all planar graphs. For every $n$-vertex graph $G\in{\mathcal{G}}$, $\operatorname{fix}(G)\geq(n/2)^{1/4}$. In other words, $\operatorname{fix}_{\mathcal{G}}(n)\geq(n/2)^{1/4}$.
@@
In contrast, applying the known bounds on the largest free sets coupled with Lemma 1 gives tight bounds for some graph families. Specifically, the linear bounds on free sets in Section 3.1 and Lemma 1 imply an $\Omega(n^{1/2})$ untangling bound for these families. Bose et al. [17] showed that there is a class of trees, namely forests of stars, ${\mathcal{S}}$, such that $\operatorname{fix}_{\mathcal{S}}\in O(n^{1/2})$. Thus the $\Omega(n^{1/2})$ untangling bound is tight for all families in Section 3.1, except for the class of cubic triconnected planar graphs. The stars in ${\mathcal{T}}$ have unbounded degree. Thus a tight bound on untangling cubic triconnected planar graphs is unknown.
@@ 4.2 Universal point subsets
A set of points $P$ is universal for a set of planar graphs if every graph from the set has a straight-line crossing-free drawing where all of its vertices map to distinct points in $P$. Arguably, the most famous open problem in graph drawing, attributed to Bojan Mohar (1988), asks if there exists a constant $c$ such that, for every $n$ there exists a pointset of size $c\cdot n$ that is universal for the class of all $n$-vertex planar graphs. It is known that for all large enough $n$, no universal pointset of size $n$ exists for the class of $n$-vertex planar graphs – as first proved by de Fraysseix et al. [34]. The same authors also proved that the $O(n)\times O(n)$ integer grid is universal for all $n$-vertex planar graphs and thus universal pointsets of size $O(n^{2})$ exist. Currently the best known lower bound on the size of a smallest universal pointset for $n$-vertex planar graphs is $1.239n-o(n)$ [52] and the best known upper bound is $n^{2}/4-O(n)$ [5]. Closing the gap between $\Omega(n)$ and $O(n^{2})$ is a major, and likely very difficult, graph drawing problem, open since 1988 [33, 34].
Various related notions have been introduced and studied in the literature with the aim of better understanding the universal pointset problem. A set $P$ of $k$ points in the plane is a universal point subset for a family of planar graphs, if for every graph $G$ in the family, there exists a straight-line crossing-free drawing in which $k$ vertices of $G$ are placed at the $k$ points in $P$. Universal point subsets were introduced by Angelini et al. [2]. They proved that a particular convex chain of $\lceil\sqrt{n}\rceil$ points is a universal point subset for the class of all $n$-vertex planar graphs. Di Giacomo et al. [37] then proved that every set of $\Omega(\log n)$ points, as well as, every set of $\lceil n^{1/3}\rceil$ points in convex position is a universal point subset for the class of $n$-vertex planar graphs.
@@ 4.3 Simultaneous Geometric Embeddings
Simultaneous geometric embeddings were introduced by Braß et al. [19]. Since then, there has been a plethora of work on the subject on many variants of the problem– see, for example, a survey by Bläsius et al. [13]. Common variants of the problem include those in which the mapping between the vertices of the two graphs is given and those in which the mapping is not given.
@@ 4.3.1 Without Mapping
A sequence of graphs $G_{1},G_{2},\dots,G_{r}$ are said to have a simultaneous geometric embedding without mapping (SGE-nomap) if there exists a pointset $P$ of size $\max\{|V(G_{1})|,\ldots,|V(G_{r})|\}$ such that each of $G_{1},\ldots,G_{r}$ has a straight-line crossing-free drawing where all of its vertices are mapped to distinct points in $P$. The following well known and still wildly open problem was asked by Braß et al. [19] in 2003, and is also listed among selected list of graph drawing problems in [18, Problem 12]: Does every pair of $n$-vertex planar graphs have a SGE-nomap (for every positive integer $n$)? The statement is known not to be true when “pair” (that is, $r=2$) is replaced by a bigger constant [21, 52], currently being $r=30$ [55]. More generally, Steiner [55] showed that for every large enough $n$ there exists a sequence of at most $(3+o(1))\log_{2}(n)$ $n$-vertex planar graphs that do not have a SGE-nomap.
While the problem of Braß et al. [19] still seems to be out of reach, in a different direction Angelini et al. [3] write: “What is the largest $k\leq n$ such that every $n$-vertex planar graph and every $k$-vertex planar graph admit a geometric simultaneous embedding without mapping? Surprisingly, we are not aware of any super-constant lower bound for the value of $k$?” The corollary below, noticed first by Dujmović [26], answered their question with the help of free sets, as detailed in the next lemma.
@@
Let $G_{1}$ be a planar graph with a free set of size at least $k$. Let $G_{2}$ be a planar graph on at most $k$ vertices. Then $G_{1}$ and $G_{2}$ admit SGE-nomap.
@@
One can combine the many lower bounds (on the size of the largest free set in a graph family) obtained in Sections 3.1, 3.2 and 3.3 to obtain results of similar flavour to 6. One example is the following: for every $n$ and every $k\leq n/4$, every $n$-vertex planar graph and every $k$-vertex triconnected cubic planar graph admits a SGE-nomap.
Note that universal point subsets of size $k$ do not imply results on SGE-nomap, akin to 6. In particular, imagine that, as in the proof of Lemma 10, one starts with a straight-line crossing-free drawing of the smaller graph $G_{2}$. Even if the larger graph $G_{1}$ is in a class of graphs that have universal point subsets of size $k$, there is no guarantee that $G_{1}$ can be drawn on the specific pointset of size $k$ used by the drawing of $G_{2}$. To be able to use universal point subsets for SGE-nomap problems, one needs the stronger variant of “all pointsets of size $k$” to be subset universal—something that is guaranteed by free sets of size $k$.
@@ 4.3.2 With Mapping
Two $n$-vertex planar graphs $G_{1}$ and $G_{2}$ on the same vertex set, $V:=V(G_{1})=V(G_{2})$, are said to have a $k$-partial simultaneous geometric embedding with mapping ($k$-PSGE-withmap) if there exists a set $V^{\prime}:=\{v_{1},\ldots,v_{k}\}\subseteq V$, and a set $P:=\{p_{1},\ldots,p_{k}\}$ of points such that each of $G_{1}$ and $G_{2}$ has a straight-line crossing-free drawing in which $v_{i}$ is mapped to $p_{i}$, for each $i\in\{1,\ldots,k\}$. PSGE-withmap on the whole vertex set (i.e. $n$-PSGE-withmap) has been widely studied, leading to mostly negative results (thus giving another motivation to introduce this partial version). For example, it is known that, for every large enough $n$, there are pairs of $n$-vertex planar graphs that do not have $n$-PSGE-withmap [19]. In fact the same is true for graphs from very simple families of planar graphs, for example: for an $n$-vertex tree and an $n$-vertex path [4], for an $n$-vertex planar graph and an $n$-vertex matching [4] and for three $n$-vertex paths [19].
The $k$-PSGE-withmap problem was introduced [^13] In [31], they use the abbreviation $k$-PSGE for what we call $k$-PSGE-withmap. Also, the definition of $k$-PSGE-withmap in [31] has one additional requirement, which states that if $v,w\in V$ are mapped to the same point in $D_{i}$ and $D_{j}$, then $v=w$. However, this additional requirement can always be met by the fact that it is possible to perturb any subset of vertices in a straight-line crossing-free drawing without introducing crossings. (This fact is used, for example, in the proof that every free-collinear set is free.) by Evans et al. [31] who proved that any two $n$-vertex trees have an $1711n/$ 17-PSGE-withmap. Their proof uses column planarity, which is the topic of the next section. Barba et al. [8] proved that any two $n$-vertex outerplanar graphs have an $n/4$-PSGE-withmap. Evans et al. [31] also observed that 4, the untangling result, implies that every pair of $n$-vertex planar graphs has an $\Omega(n^{1/4})$-PSGE-withmap. Namely, start with a straight-line crossing-free drawing of $G_{1}$. Since $G_{1}$ and $G_{2}$ have the same vertex set, the drawing of $G_{1}$ (or rather the positions of its vertices in the plane) defines a straight-line drawing of $G_{2}$ (that almost certainly has crossings). Untangling $G_{2}$ while keeping $\Omega(n^{1/4})$ of its vertices fixed (which is possible by 4) gives the result.
@@
However, the above untangling argument fails if we try to apply it one more time. The following generalization of the $k$-PSGE problem to more than two graphs illustrates this: Given any set $Q:=\{G_{1},\ldots,G_{r}\}$ of planar graphs on the same vertex set, $V$, we say that the graphs in $Q$ have a $k$-partial simultaneous geometric embedding with mapping ($k$-PSGE-withmap) if there exists a set $V^{\prime}:=\{v_{1},\ldots,v_{k}\}\subseteq V$, and a set $P:=\{p_{1},\ldots,p_{k}\}$ of points such that each graph in $Q$ has a straight-line crossing-free drawing in which $v_{i}$ is mapped to $p_{i}$, for each $i\in\{1,\ldots,k\}$.
@@
We start by presenting a lemma about PSGE-withmap for two graphs via free sets. Its proof uses the common trick of taking advantage of the fact that points in the plane have two degrees of freedom, and thus one ordering can be imposed on the ${x}$-coordinates and a different ordering on ${y}$-coordinates (see for example [7]).
@@
6 and 11 gives another proof of 7. By 6 (applied to $G_{1}$ and $V$), there exists $S_{1}\subseteq V$ of size at least $\sqrt{n/2}$ that is a free set in $G_{1}$. By 6 (applied to $G_{2}$ and $S_{1}$), there exists $S\subseteq S_{1}$ of size at least $\sqrt{|S_{1}|/2}\geq n^{1/4}/2^{3/4}$. Thus $S$ is a free set in $G_{1}$ and in $G_{2}$. By Lemma 11, $G_{1}$ and $G_{2}$ have a $\lceil n^{1/4}/2^{3/4}\rceil$-PSGE-withmap. In exactly the same way, Lemmata 3 and 11 give a proof that any two $n$-vertex graphs, each of which is one of the following: a tree, an outerplanar graph, Halin graph, or a square graph, have a $(n/4)$-PSGE-withmap.
@@
Now apply Lemma 11 to the graphs $G_{1}$ and $G_{2}$ with the set $S$ to obtain straight-line crossing-free drawing of both graphs in which each vertex in $S$ (and therefore, the vertices in $S^{\prime}$) appears at the same location in both drawings. In these drawings, the ${y}$-coordinates of the vertices in $S^{\prime}$ are increasing. By the definition of free set (exchanging the roles of ${x}$- and ${y}$-coordinates), each of $G_{3},\ldots,G_{k}$ has a straight-line crossing-free drawing  in which each vertex in $S^{\prime}$ has the same location in each drawing. ∎
@@
If the graphs in $Q$ each come from one of the classes covered by 1, then repeated applications of 1 gives a set $S$ of size at least $n/2^{r}$ that is free in each of $G_{1},\ldots,G_{r}$, in which case the bound in 7 becomes $n^{1/2^{r-2}}/4$.
@@ 4.4 Column planarity
Given a planar graph $G$, a set $R\subseteq V(G)$ is column planar in $G$ if the vertices of $R$ can be assigned distinct ${x}$-coordinates such that for an assignment of ${y}$-coordinates to the vertices in $R$ such that the resulting $|R|$ points have no three on a line, there exists a straight-line crossing-free drawing of $G$ in which each vertex $v\in R$ is placed to its assigned coordinates. $R$ is strongly column planar if the “no three on a line condition” is removed from the definition. The following lemma is immediate from these definition and 1:
@@
A (sub)set of vertices of a planar graph is a strongly column planar set if and only if it is an unordered free set.
@@
Column planar sets were introduced by Evans et al. [31] motivated by applications to partial simultaneous geometric embeddings. Notions similar to column planarity were studied by Estrella-Balderrama et al. [30] and Giacomo et al. [38]. Barba et al. [7] proved that $n$-vertex trees have column planar sets of size $1714n/$ 17. Note that that does not imply that trees have free sets of that size, since being column planar does not imply being strongly column planar. In particular, if $G$ is a 3-cycle, then $V(G)$ is a column planar set for any assignment of ${x}$-coordinates to $V(G)$. However, no ${x}$-coordinate assignment can be turned into a straight-line crossing-free drawing with all the vertices having the same ${y}$-coordinate. Thus $V(G)$ is not strongly column planar.